OpenStudy (anonymous):

prove that (2^(2n)+2(^n+1)+1)^(1/2)-(2^(2n)-2(^n+1)+1)^(1/2)=2 plz help :)

hartnn (hartnn):

can you expand \((2^n+1)^2=...?\) \using the formula \((a+b)^2=a^2+2ab+b^2\)

OpenStudy (anonymous):

yes thanx a lot

hartnn (hartnn):

so, could you prove that ?

hartnn (hartnn):

you'll need ,\((2^n-1)^2=...?\) also, then just subtract those 2 results...